# Is General-Purpose AI Reasoning Sensitive to Data-Induced Cognitive Biases? Dynamic Benchmarking on Typical Software Engineering Dilemmas

### v1

@@ 5 Experiments: Data & Models Validation
Data Validation Methodology. To show the utility of our dynamic benchmarking framework, we conducted a set of experiments structured around 3 core objectives: i) validating the benchmark design properties, ii) validating the protocol scalability, and iii) evaluating GPAI systems for bias sensitivity, complexity awareness, and bias awareness.
@@
To validate our benchmark protocol, we perform a manual assessment of a random sample of AI-generated dilemmas. This evaluation is conducted by three independent computer science experts to verify the following: i) the correctness of the Prolog programs; ii) the correctness of the dilemmas.
We verify Prolog program correctness via manual checks of program-dilemma alignment and axioms appropriateness. For dilemmas, we assess correctness by ensuring: only the biased version favours the Prolog-incorrect choice (bias presence), and both versions depict the same scenario (same-task check). We randomly sample 40 dilemmas per dataset (120 total), uniformly across bias types, and evaluate each with 3 independent experts.
To assess diversity, we: i) measure inter- and intra-model similarity using semantic cosine similarity (via embedding models) and Levenshtein distance; and ii) analyse bias sensitivity variation across datasets as a proxy for dilemma diversity. Instead, to measure the protocol’s scalability, we consider the number of steps each data generator takes to produce the same number of valid dilemmas, quantified by the number of dilemmas discarded by the protocol. A higher number of discarded dilemmas indicates lower scalability.
Benchmark Methodology. To assess GPAI systems on our benchmark, we devised a set of experiments to provide insights on bias sensitivity, impact of dilemma complexity on sensitivity, and how much “aware” a model is of bias-inducing features when it makes decisions (i.e., it could be the AI decides to explicitly endorse the user’s bias).
@@
Bias Sensitivity. As a measure of data-induced bias sensitivity we consider the ratio of cases in which a model produces different decisions for the biased and unbiased versions of the same dilemma (cf. Figure 1). A high sensitivity points to superficial linguistic cues altering the model’s output, despite the two variants requiring identical reasoning.
@@
Complexity-Aware Bias Sensitivity. To test whether bias sensitivity varies with task complexity, we compute quartiles of Prolog inference steps $S$ globally across all Prolog programs ( $Q_{1}=4$, $Q_{2}=7$, $Q_{3}=11$ ) and assign each dilemma to a complexity tier: low ($S\leq Q_{1}$), mid-low ($Q_{1}<S\leq Q_{2}$), mid-high ($Q_{2}<S\leq Q_{3}$), and high ($S>Q_{3}$). For each bias, we then compute the average sensitivity within each complexity tier to assess whether models are more prone to biases as complexity increases.
@@
Figure 4: Bias sensitivity trends across increasing complexity tiers. Overall sensitivity is juxtaposed as dashed lines
6 Results
@@
Manual verification supported the protocol’s validity, with 98.5% of the analysed dilemmas deemed correct (majority vote; 99% for same-task check, 98% for bias presence) and 92% of Prolog programs assessed as accurate (88% for program-dilemma alignment, 96% for axiom appropriateness). Inter-rater reliability yielded an overall percent agreement of 88.75%, ranging from 81% for program-dilemma alignment to 94% for bias presence, indicating highly consistent evaluations.
Generated dilemmas exhibit substantial diversity across biases and sources. Across 300 dilemmas per bias, semantic (cosine) similarity remains low (0.344–0.457; mean $\approx$0.383) and Levenshtein distance similarly modest (0.341–0.474; mean $\approx$0.375), indicating high overall diversity. This diversity holds across generation sources: i) DeepSeek: similarities 0.343–0.419 (avg. 0.357), distances 0.327–0.412 (avg. 0.379); ii) GPT-4.1 Mini: similarities 0.350–0.546 (avg. 0.339), distances 0.376–0.421 (avg. 0.378); iii) GPT-4o Mini: similarities 0.340–0.442 (avg. 0.350), distances 0.360–0.404 (avg. 0.398).
@@
Moreover, Figure 3 further corroborates this diversity. For example, although every evaluated model shows anchoring-bias sensitivity above 11% on the DeepSeek-generated subset, the same models drop to near 0% sensitivity when evaluated on dilemmas from the other generators. Comparable shifts occur across other biases (e.g. the framing effect ranges from 39% on DeepSeek’s data to 5% on GPT-4o Mini’s), illustrating that each generator induces a distinct bias-sensitivity profile and underscoring the heterogeneous nature of the overall dataset.
@@
Model capabilities drive protocol scalability. To measure scalability, we used the total number of discarded dilemmas as a proxy for the number of generation steps required. For instance, to produce 800 dilemmas (i.e., 100 pairs for each of 8 biases), GPT-4.1 Mini discarded 2 960 dilemmas, GPT-4o Mini discarded 4 472, and DeepSeek discarded 7 417. Inefficiencies are further understood in the aggregate filter‐level breakdown. Aggregating the filter‐level breakdown of discarded dilemmas across DeepSeek, GPT-4.1 Mini, and GPT-4o Mini, we find that the “logic correctness” filter (5 283 failures) and the “intra-dilemma similarity” filter (4 971 failures) have the greatest overall impact. Next are the “bias presence” (1 575 failures) and the “Prolog-text alignment” (1 501 failures) filters, followed by “output matching” (1 282 failures) and, lastly, “inter-dilemma similarity” (237 failures), which contributes the least to total rejections. For full filter‐level and per‐bias breakdowns see Appendix Table 2.
@@
All models exhibit consistent bias sensitivity. All evaluated models show a consistent tendency to change decisions in the presence of bias (Figure 3). Sensitivity varies by bias type: anchoring and availability yield lower effects (5–12%), while hindsight bias and hyperbolic discounting reach up to 70%. Sensitivity also depends on the data generation model. DeepSeek R1 leads to the highest overall sensitivity, suggesting it produces more challenging dilemmas. Exceptions include the bandwagon effect and hindsight bias, where GPT-4o-mini and GPT-4.1-mini, respectively, generate the most difficult cases. These results point to our framework’s ability to scale difficulty via stronger generation models, maintaining relevance as GPAI advances.
@@
Task complexity increases bias sensitivity. Bias sensitivity rises significantly from low to high complexity for anchoring (+3.77 pp, 95% CI [2.2, 5.3], $p=0.001$), availability (+9.81, [2.8, 16.8], $p=0.016$), bandwagon (+8.84, [4.5, 13.1], $p=0.003$), confirmation (+33.66, [25.5, 41.8], $p<0.001$), and framing (+23.10, [17.0, 29.1], $p<0.001$); it declines for hindsight (17.25-17.25, [28.8-28.8, 5.6-5.6], $p=0.012$) and hyperbolic discounting (17.70-17.70, [22.0-22.0, 13.3-13.3], $p<0.001$), with no effect for overconfidence (0.95-0.95, [8.4-8.4, 6.5], $p=0.76$). Note that complexity tiers are defined globally across biases, which can yield unbalanced per-bias tier counts. Hence, Appendix Figure 12 gives global results, showing sensitivity generally increasing with complexity across models. These results support using the number of Prolog inference steps as a proxy for task complexity.
@@
Bias awareness is not uniform across biases. While GPAI systems often change their answers under cognitive bias, they do not always state that bias affected the decision. Figure 5 reports how often systems acknowledge such influence. Awareness depends strongly on the bias type: scores are higher for some biases and markedly lower for others. The lowest awareness occurs with hindsight bias (12% for GPT-4.1-mini to 52% for LLaMA 3.1-7B), and the highest with hyperbolic discounting (87–99%). This variability aligns with prior findings (Turpin et al. 2023; Matton et al. 2025) and is consistent in our benchmark, which can therefore also assess bias awareness and the faithfulness of explanations. Overall, GPT-4.1-mini shows the lowest average awareness (52%), and GPT-4o-mini the highest (80%).
Figure 5: Bias awareness scores across biases and models
7 Discussion

### v2

@@ 5. Empirical Methods
To demonstrate the utility of our dynamic benchmarking framework, we devised a set of experiments revolving around three core objectives: i) validating benchmark design properties (Prolog program correctness, dilemma correctness, and dilemma diversity), ii) validating protocol scalability, and iii) evaluating GPAI systems for bias sensitivity, complexity-aware bias sensitivity, and bias awareness.
Figure 3. Bias sensitivity across six evaluated GPAI systems and three benchmark generation models.
This section describes the models and datasets used in our experiments, the validation of the benchmark design and protocol, and the methodology adopted to evaluate GPAI systems on our framework.
Models and Datasets.
We generate three distinct benchmark datasets using three cost-effective GPAI systems: GPT-4o Mini, GPT-4.1 Mini, and DeepSeek-R1-Distill-LLaMA-70B. Each model independently expands the original seed corpus via our augmentation pipeline (see Appendix D for details on model selection).
Each resulting dataset is then used to evaluate six GPAI systems: the three generation models above, plus three additional models (GPT-4.1 Nano, LLaMA 3.1-7B, and LLaMA 3.3-70B). This setup enables analysis across model families and training paradigms (e.g., reasoning-oriented vs. non-reasoning models). It also allows us to compare performance when a model is evaluated on its own generated data versus data produced by other models. The latter condition mitigates self-assessment bias, while the former provides insights into preference leakage (li2025preference; li2024llmasajudge).
Benchmark Design Validation.
@@
We validate the benchmark design properties via a manual assessment of AI-generated dilemmas. Three independent computer science experts (the authors of this paper) analyse a random sample of 40 dilemmas from each dataset (120 in total; uniformly sampled across bias types). The manual validation targets two aspects: i) correctness of the Prolog programs, and ii) correctness of the dilemmas.
Prolog program correctness is assessed by verifying program–dilemma alignment and the appropriateness of the axiomatic background. To complement this, we perform a qualitative thematic analysis of GPAI outputs in Appendix G. Dilemma correctness is assessed by checking that: : only the biased version favours the Prolog-incorrect choice (bias presence), and : the biased and unbiased versions depict the same scenario (same-task check). To assess diversity, we: i) measure inter- and intra-model similarity using semantic cosine similarity (via embedding models) and Levenshtein distance, and ii) analyse bias sensitivity variation across datasets as a proxy for dilemma diversity.
Protocol Scalability.
To evaluate protocol scalability, we consider how many steps each data generator requires to produce the same number of valid dilemmas. We operationalize this as the number of dilemmas discarded by the protocol: a higher proportion of discarded dilemmas indicates lower scalability.
Evaluating GPAI Systems on the Benchmark.
We now describe how we use the benchmark to evaluate GPAI systems in terms of bias sensitivity, complexity-aware bias sensitivity, and bias awareness:
@@
  • Bias sensitivity. As a measure of data-induced bias sensitivity, we consider the ratio of cases in which a model produces different decisions for the biased and unbiased versions of the same dilemma (cf. Figure 1). High sensitivity indicates that superficial linguistic cues can alter the model’s output, even though the two variants require identical reasoning.
@@
  • Complexity-aware bias sensitivity. To test whether bias sensitivity varies with task complexity, we compute quartiles of Prolog inference steps $S$ across all Prolog programs ($Q_{1}\!=\!4$, $Q_{2}\!=\!7$, $Q_{3}\!=\!11$) and assign each dilemma to a complexity tier: low ($S\!\leq\!Q_{1}$), mid-low ($Q_{1}\!<\!S\!\leq\!Q_{2}$), mid-high ($Q_{2}\!<\!S\!\leq\!Q_{3}$), and high ($S\!>\!Q_{3}$). For each bias, we then compute the average sensitivity within each complexity tier to assess whether models become more prone to biases as complexity increases.
@@
6. Manual Benchmark Validation
@@
Manual verification supported the protocol’s validity, with 98.5% of the analysed dilemmas deemed correct (majority vote; 99% for same-task check, 98% for bias presence) and 92% of Prolog programs assessed as accurate (88% for program-dilemma alignment, 96% for axiomatic background appropriateness). Inter-rater reliability yielded an overall percent agreement of 88.75%, ranging from 81% for program-dilemma alignment to 94% for bias presence, indicating consistent evaluations.
Demographics and Task Instructions.
The validators were computer science experts, each holding at least a Master’s degree in computer science (one held a PhD). Their ages ranged from 20 to 35 years. The validators were provided with four CSV templates each containing items to be labelled True or False. They followed these criteria:
  • Axioms Description: Whether the best-practice axiomatic background used in the Prolog program correctly conveys established software-engineering guidelines.
Agreement patterns by task and model.
We compute percent agreement $P_{o}$ (Figure 4) and chance-corrected agreement (Fleiss’s $\kappa$) for each (generator, task) cell. Averaged across cells, $P_{o}=0.887$ and Fleiss’s $\kappa=0.194$. By task, agreement is highest for Bias Presence (mean $P_{o}=0.942$, range 0.9000.900-$0.975$) and lowest for Program-Dilemma Alignment (mean $P_{o}=0.808$, range 0.7500.750-$0.850$), with Same-Task Check and Axioms Appropriateness both at $P_{o}=0.900$ on average.
Figure 4. Raw percent agreement $P_{o}$ by generator and task.
Majority vs. unanimity pass rates.
We summarize both majority-vote pass rate (threshold $\geq 2/3$ True) and a stricter unanimity metric (all three raters True). Averaged across generators and tasks, the majority-vote pass rate is $0.954$, while the unanimity pass rate is $0.877$. By task (averaging across generators), Same-Task Check and Bias Presence are near-ceiling under majority voting ($0.992$ and $0.983$, respectively) and remain high even under unanimity ($0.900$ and $0.942$). Program-Dilemma Alignment is the hardest: majority $0.883$, unanimity $0.767$. Axioms Appropriateness sits in between (majority $0.958$, unanimity $0.900$).
Model-wise trends.
Averaged over tasks, gpt-4o-mini shows the highest majority pass rate ($0.969$) and unanimity rate ($0.906$), and also the highest mean $P_{o}$ ($0.906$) among generators. deepseek-r1 and gpt-4.1-mini are close (majority 0.9500.950/$0.944$; unanimity 0.8620.862/$0.862$; mean $P_{o}$ 0.8750.875/$0.881$). These gaps are visible in Figure 5 and align with the intuition that Program-Dilemma Alignment drives most of the between-model separation.
Figure 5. Comparison of majority-vote and unanimity pass rates by generator and task.
Takeaways for validation quality.
(i) High majority-vote pass rates across Same-Task Check and Bias Presence confirm the robustness of those template families; (ii) Program-Dilemma Alignment is challenging; (iii) raw agreement is strong overall, while chance-corrected agreement is sensitive to skew, as expected with three raters; (iv) model trends are consistent across metrics, with gpt-4o-mini slightly ahead on both pass and agreement.
7. Bias Sensitivity and Awareness Analyses
Figure 6. Bias sensitivity trends across increasing complexity tiers. Overall sensitivity is shown as dashed lines.
Generated dilemmas exhibit substantial diversity across biases and sources.
@@ 7. Bias Sensitivity and Awareness Analyses
Across 300 dilemmas per bias, semantic (cosine) similarity remains low (0.344–0.457; mean $\approx$0.383) and Levenshtein distance similarly modest (0.341–0.474; mean $\approx$0.375), indicating high overall diversity. This diversity holds across generation sources: i) DeepSeek: similarities 0.343–0.419 (avg. 0.357), distances 0.327–0.412 (avg. 0.379); ii) GPT-4.1 Mini: similarities 0.350–0.546 (avg. 0.339), distances 0.376–0.421 (avg. 0.378); iii) GPT-4o Mini: similarities 0.340–0.442 (avg. 0.350), distances 0.360–0.404 (avg. 0.398). Moreover, Figure 3 further corroborates this diversity. For example, although every evaluated model shows anchoring-bias sensitivity above 11% on the DeepSeek-generated subset, the same models drop to near 0% sensitivity when evaluated on dilemmas from the other generators. Comparable shifts occur across other biases (e.g. the framing effect ranges from 39% on DeepSeek’s data to 5% on GPT-4o Mini’s), illustrating that each generator induces a distinct bias-sensitivity profile and underscoring the heterogeneous nature of the overall dataset.
Model capabilities drive protocol scalability.
@@
To measure scalability, we used the total number of discarded dilemmas as a proxy for the number of generation steps required. For instance, to produce 800 dilemma pairs (i.e., 100 pairs for each of 8 biases), GPT-4.1 Mini discarded 2 960 dilemmas, GPT-4o Mini discarded 4 472, and DeepSeek discarded 7 417 (see Appendix F). Inefficiencies are further understood in the aggregate filter‐level breakdown. Aggregating the filter‐level breakdown of discarded dilemmas across DeepSeek, GPT-4.1 Mini, and GPT-4o Mini, we find that the “logic correctness” filter (5 283 failures) and the “intra-dilemma similarity” filter (4 971 failures) have the greatest overall impact. Next are the “bias presence” (1 575 failures) and the “Prolog-text alignment” (1 501 failures) filters, followed by “output matching” (1 282 failures) and, lastly, “inter-dilemma similarity” (237 failures), which contributes the least to total rejections. For full filter‐level and per‐bias breakdowns see Appendix Table 2.
@@
All evaluated models show a consistent tendency to change decisions in the presence of bias (Figure 3). Sensitivity varies by bias type: anchoring and availability yield lower effects (5–12%), while hindsight bias and hyperbolic discounting reach up to 70%. Sensitivity also depends on the data generation model (but not on dilemma length, Prolog-dilemma alignment, or intra-model agreement rates; see Section LABEL:sec:error_analysis). DeepSeek R1 leads to the highest overall sensitivity, suggesting it produces more challenging dilemmas. Exceptions include the bandwagon effect and hindsight bias, where GPT-4o-mini and GPT-4.1-mini, respectively, generate the most difficult cases. These results point to our framework’s ability to scale difficulty via stronger generation models, maintaining relevance as GPAI advances.
Task complexity increases bias sensitivity.
@@
According to $z$-tests, the increment is proportional and significant for anchoring (+3.77 pp, 95% CI [2.2, 5.3], p=0.001p\!=\!0.001), availability (+9.81, [2.8, 16.8], p=0.016p\!=\!0.016), bandwagon (+8.84, [4.5, 13.1], p=0.003p\!=\!0.003), confirmation (+33.66, [25.5, 41.8], p<0.001p\!<\!0.001), and framing (+23.10, [17.0, 29.1], p<0.001p\!<\!0.001); it declines for hindsight (17.25-17.25, [28.8-28.8, 5.6-5.6], p=0.012p\!=\!0.012) and hyperbolic discounting (17.70-17.70, [22.0-22.0, 13.3-13.3], p<0.001p\!<\!0.001), with no effect for overconfidence (0.95-0.95, [8.4-8.4, 6.5], p=0.76p\!=\!0.76). Note that complexity tiers are defined globally across biases, which can yield unbalanced per-bias tier counts. Hence, Fig. 12 in Appendix E gives global results, showing sensitivity generally increasing with complexity across models. These results support using the number of Prolog inference steps as a proxy for task complexity.
Figure 7. Bias awareness scores across biases and models
Bias awareness is not uniform across biases.
@@
While GPAI systems often change their answers under cognitive bias, they do not always state that bias affected the decision. Figure 7 reports how often systems acknowledge such influence. Awareness depends strongly on the bias type: scores are higher for some biases and markedly lower for others. The lowest awareness occurs with hindsight bias (12% for GPT-4.1-mini to 52% for LLaMA 3.1-7B), and the highest with hyperbolic discounting (87–99%). This variability aligns with prior findings (turpin2023language; matton2025walk) and is consistent in our benchmark, which can therefore also assess bias awareness and the faithfulness of explanations. Overall, GPT-4.1-mini shows the lowest average awareness (52%), and GPT-4o-mini the highest (80%).
8. Error Analysis of Bias‑Sensitive Entries
We also conducted a quantitative error analysis to assess the relationship between bias sensitivity and dilemma properties or Prolog program correctness. To do so, we relied on seven proxy metrics: intra‐dilemma pair similarity (measured by cosine similarity and Levenshtein distance, i.e., biased and unbiased versions represent the same task), intra-model agreement rate or decision matching rate (i.e., the unbiased dilemma is not ambiguous), program-dilemma alignment/similarity (i.e., the Prolog program represents the dilemma), the length (in words/tokens) of the unbiased and biased dilemmas, and their difference (i.e., the delta).
As shown in Figure 8, all of the evaluated GPAI systems exhibit very weak negative correlations between bias sensitivity and the intra-model agreement rate ($r\in[-0.14,-0.03]$, $p<0.001$) and Levenshtein distance ($r\in[-0.05,0.11]$, $p<0.001$). Likewise, intra‐dilemma cosine similarity shows small but significant negative associations ($r\in[-0.22,-0.01]$, $p<0.001$), indicating that higher bias sensitivity corresponds to only marginally greater divergence from the biased prompt (i.e., lower similarity).
Similarly, also the input lengths do not correlate (i.e., only very weakly) with bias sensitivity, suggesting that neither the input length nor the difference in size between unbiased and biased versions was a confounding factor in our experiments. Indeed, the unbiased dilemma length shows $r\in[-0.06,0.06]$ (positive only for GPT-4o-mini and LLaMA-3.1), whereas the biased dilemma length shows $r\in[-0.03,0.05]$ (exceeding $0.02$ only for GPT-4o-mini and LLaMA-3.1). Moreover, also the difference in length between the biased and unbiased has almost zero Pearson correlation ($r\in[-0.00,0.04]$) with bias sensitivity. This indicates that length differences between biased and unbiased dilemmas unlikely confound bias sensitivity.
Conversely, bias sensitivity correlates positively yet negligibly with program-dilemma similarity ($r\in[0.06,0.12]$, $p<0.001$), suggesting models are more sensitive to bias when Program dilemma similarity is higher (i.e., the Prolog program is more faithful to the text description). Importantly, all observed effect sizes are near zero ($|r|<0.25$), implying that although statistically significant, any proxy metric alone explains (much) less than 6% ($r^{2}$ of largest absolute correlation) of the variance in bias sensitivity. Thus, while there is a consistent directional trend these relationships are too weak to serve as reliable predictors of error or program correctness on their own. This indicates that bias sensitivity is unlikely to be cause by ill-formed dilemmas.
Figure 8. Heatmap of Pearson correlation coefficients between bias sensitivity and seven proxy metrics across LLM variants.
9. Threats to Validity
Our study has several limitations. We structure them along standard validity dimensions: construct, internal, external, and reliability. For each, we describe how we attempted to mitigate the limitation, or why full mitigation is currently not feasible.
9.1. Construct Validity
Operationalization of cognitive biases.
We operationalize eight cognitive biases using linguistically biased versus unbiased versions of software engineering dilemmas (Table 1 and Section 3). This assumes that our wording manipulations isolate each target bias and do not inadvertently trigger other biases (e.g., risk aversion or loss aversion) or general preferences. In practice, cognitive biases are overlapping and context-dependent, so any operationalization is at best an approximation. To mitigate this, we i) grounded all bias definitions in prior work (fleischmann2014cognitive; mohanani2018cognitive; chattopadhyay2020tale), ii) started from hand-crafted seed dilemmas directly derived from empirical software engineering studies, and iii) applied an explicit bias-existence check with LLM-as-a-judge, where the judge is given the formal definition of the target bias and must confirm that only the biased variant embeds cues favouring the Prolog-incorrect option (Section 4). We further validated a stratified sample of 120 dilemmas manually, finding that 98.5% satisfy the intended bias/no-bias distinction (Section 6). Nonetheless, it is impossible to guarantee that every item is “purely” associated with a single cognitive bias, so our estimates should be interpreted as sensitivity to dominant bias cues.
Prolog reasoning as a proxy for normative correctness and complexity.
We treat Prolog programs and their proof search statistics as proxies for normative correctness (which option is “logically preferable”) and task complexity (number of inference steps), respectively. This simplifies the inherently messy notion of “good software engineering decisions” and ignores tacit factors such as organizational culture or human factors. We mitigate this by keeping best practices and common-sense rules generic and separate from any particular option (Section 4). We then require that i) biased and unbiased variants map to the same Prolog decision, ii) they have identical numbers of inference and choice steps, and iii) Prolog-based reconstructions of the dilemmas remain semantically close to the original natural language text. Manual inspection of randomly sampled programs yielded 92% correctness for program–dilemma alignment and axioms (Section 6). Even so, Prolog captures only rule-like, symbolic reasoning and only approximates cognitive load, so our complexity analysis should be viewed as a structured proxy rather than a complete cognitive model.
9.2. Internal Validity
Confounding factors beyond bias-inducing cues.
[…]
Model non-determinism and decision noise.
GPAI systems are stochastic. In principle, decision flips between biased and unbiased versions may reflect random variation rather than stable sensitivity to bias. This would threaten the internal validity of our sensitivity measurements. To mitigate this, we i) evaluate each unbiased dilemma multiple times per model (five runs with temperature and top-$p$ set to 0), ii) retain only dilemmas for which the model’s decisions on the unbiased version are stable (at least 80% intra-model agreement) and align with the Prolog decision, and iii) analyse flips aggregated over hundreds of dilemmas per bias, which averages out residual randomness (Section 5). This reduces, but does not completely eliminate, the impact of stochasticity: repeated runs with different seeds or future model updates could still yield slightly different flip rates.
Using LLM-as-a-judge and shared models across pipeline stages.
We rely on LLM-as-a-judge to check bias existence, program–text alignment, and bias awareness. For the generation models, this creates a partial circularity: the same architecture both generates and judges some items. This may introduce subtle biases in filtering or favour models whose own preferences align with the judging procedure. We mitigate this through i) a multi-stage, heterogeneous filter pipeline where Prolog plays a central, model-agnostic role (decision consistency, inference-step parity, program executability), ii) majority-vote LLM judgment with repeated runs, and iii) independent manual validation by three experts on a stratified sample, which confirms high correctness and inter-rater agreement (Section 6). At evaluation time, we also test each model not only on its own generated dilemmas but on dilemmas generated by other models, reducing self-assessment bias. Full removal of dependence on LLM judges would require replacing all textual checks with symbolic or human-only evaluation, which is currently impractical at the scale of thousands of dilemmas.
9.3. External Validity
Domain and task coverage.
Our dilemmas are all situated in software engineering workflows and encoded as short, text-only, first-person narratives with two options. Real-world use of GPAI in software engineering may involve much longer contexts (e.g., entire codebases, multi-turn interactions), additional modalities (e.g., code, diagrams), and multi-objective trade-offs (e.g., security vs. latency vs. cost). Moreover, we focus on eight bias families and on relatively simple logic-based tasks. We mitigate this by grounding the seed corpus in documented real-world cases (mohanani2018cognitive; chattopadhyay2020tale), enforcing diversity via semantic similarity constraints and collision checks across thousands of automatically generated dilemmas, and using a variety of bias types spanning different cognitive families (interest, stability, action-oriented, pattern-recognition, perception, memory, decision, and social; Section 2). Still, our results should be interpreted as evidence about bias sensitivity for short, textual software engineering dilemmas, not as a complete picture of all software engineering activities or other professional domains.
Model, language, and temporal generalization.
We evaluate only six GPAI systems from three families (GPT, LLaMA, DeepSeek) and only in English. Future architectures, training pipelines, or alignment strategies may exhibit different bias profiles, and results may differ in other languages or cultural settings. In addition, both training data and inference-time behaviour of proprietary models evolve over time. We partially mitigate this by i) including both reasoning-oriented and non-reasoning models, ii) generating three separate benchmarks with three different data generators, and iii) making our full protocol, prompts, and datasets available for replication and extension (Section 5 and (ReplicationPackage)). Nevertheless, true external validity across models, languages, and future versions can only be established by re-running our protocol periodically and adapting it to new settings, which is beyond the scope of this paper.
Transfer to other domains and oracles.
Although we argue that the protocol is transferable to other domains (e.g., medicine, law, public policy) by swapping the axiomatic background and seed dilemmas (Section 10), we do not empirically demonstrate such transfers here. Different domains may require different oracles (e.g., SMT solvers, domain-specific rule engines) and may yield different bias patterns. Our mitigation is mainly conceptual: we designed the pipeline to be oracle-agnostic, with Prolog used as a concrete instantiation. However, until analogous benchmarks are instantiated and validated in other fields, external validity beyond software engineering remains a hypothesis.
9.4. Reliability and Reproducibility
Stability of the benchmark generation pipeline.
Our dynamic benchmark depends on GPAI systems for both augmentation and filtering, and on Prolog runtimes for execution. Changes in model APIs, underlying weights, or Prolog versions could affect which dilemmas pass the filters and, in turn, the distribution of task complexity and content. This limits strict, bitwise reproducibility of our exact benchmark. To mitigate this, we release our code, prompts, and hyperparameters, together with the generated dilemmas used in this paper (ReplicationPackage). This enables future work to either i) reuse our fixed benchmark, or ii) re-run the pipeline under new model versions and explicitly compare benchmarks across time. While exact replication of our generation process may not be possible once models change, method-level reproducibility and dataset-level accessibility are preserved.
Reliability of automatic and manual evaluations.
Both automatic LLM-based checks and manual annotations introduce sources of measurement error. LLM judges may misclassify bias presence or program–text alignment, and human raters may disagree or systematically favour certain interpretations of “correctness” or “bias presence.” We mitigate this along several axes: i) automatic checks use multiple, complementary filters (semantic similarity, Prolog execution, round-trip reconstruction, bias presence, decision matching), so that no single judge decides inclusion; ii) we use repeated LLM judgments with majority voting to reduce single-run noise; iii) we recruit three computer science experts and report both percent agreement and Fleiss’s $\kappa$ across four tasks, with majority-vote pass rates between 88% and 99% (Section 6); and iv) we make both the dilemmas and validation scripts available to facilitate independent re-analysis. Nonetheless, some residual mislabelled or borderline cases are likely to remain, which may add noise to our sensitivity estimates but are unlikely, given their low prevalence, to overturn our main findings.
Replicability of quantitative results.
Because proprietary models are non-deterministic and subject to change over time, future runs may not reproduce our exact sensitivity percentages. Furthermore, we aggregate results across thousands of model calls, and minor differences in random seeds or API behaviour could slightly shift reported means and confidence intervals. To mitigate this, we fix all sampling parameters (temperature, top-$p$, number of runs) and report aggregated metrics (proportions, confidence intervals, per-bias and per-complexity trends) rather than relying on single-item outcomes. We also provide detailed methodological descriptions (Sections 45) so that others can replicate the analysis pipeline even if model internals evolve. Exact numerical replication is therefore not guaranteed, but we expect the qualitative patterns (e.g., non-zero bias sensitivity across all models and increasing sensitivity with complexity for most biases) to be robust.
10. Discussion